# Reducing Size Bias in Sampling Infectious Disease Spread on Networks

### v2

@@ Abstract
Epidemiological models can aid policymakers in reducing disease spread by predicting outcomes based on disease dynamics and contact network characteristics. Calibrating these models requires representative network samples. In this connection, we investigate two sampling algorithms, Random Walk (RW) and Metropolis-Hastings Random Walk (MHRW), across three network types: Erdős-Rényi (ER), Small-World (SW), and Scale-Free (SF). Disease transmission is simulated using a stochastic susceptible-infected-recovered (SIR) framework. Our findings show that RW overestimates infected individuals and secondary infections by $25\%$ for ER and SW networks due to size bias, favouring highly connected nodes. MHRW, which corrects for size bias, provides estimates that are more consistent with the underlying network. Also, both methods yield estimates closer to the underlying network for time-to-infection. However, sampling SF networks exhibits significant variability, for both algorithms. Removing duplicate sampled nodes reduces MHRW’s accuracy across all network types. We apply both algorithms to a cattle movement network of 46,512 farms, exhibiting ER, SW, and SF network features. RW overestimates infected farms by approximately $100\%$ and secondary infections by $>900\%$, reflecting size bias whereas MHRW estimates align closely with the cattle network dynamics. Time-to-infection estimates reveal that RW underestimates by approximately $40\%$, while MHRW slightly overestimates by $10\%$. Estimates differ greatly when duplicate nodes are removed. These findings underscore choosing algorithms based on network structure and disease severity. RW’s conservative estimates suit high-mortality, fast-spreading diseases, while MHRW provides precise interventions suitable for less severe outbreaks. These insights can guide policymakers in optimising resource allocation and disease control strategies.
@@
networks , sampling , disease modelling , size bias , cattle network , policymaking , SIR model
MSC:
92D30 , 05C82 , 62P10 , 91D30 , 90B15
@@
Assessed RW and MHRW sampling for disease metrics in ER, SW, and SF networks.
MHRW outperformed RW in ER and SW, reducing size bias for accurate estimates.
SF networks showed high variability, suggesting alternative sampling methods.
RW sampling overestimated disease metrics, while MHRW yielded precise results.
@@ 1 Introduction
During an epidemic, accurately estimating metrics such as the number of infected individuals, effective reproduction number, and time-to-infection is crucial for effective mitigation (Nunes et al., 2024). Mathematical modelling provides a suite of tools, frameworks, and methodologies to generate these estimations (Vynnycky and White, 2010). However, estimating disease metrics on networks is challenging due to factors such as network complexity (Avraam and Hadjichrysanthou, 2025) and the influence of the chosen data sampling method (Danon et al., 2011). The Random Walk (RW) algorithm, a commonly used sampling method, is prone to size bias, leading to over-representation of highly connected nodes and potentially skewed estimates of disease spread. Metropolis-Hastings Random Walk (MHRW) addresses this limitation by correcting for size bias, offering a more balanced representation of network nodes.
In this paper, we statistically compare RW and MHRW to evaluate their performance in estimating disease metrics. Using stochastic susceptible-infected-recovered (SIR) simulations, we assess these algorithms across three network types, Erdős-Rényi (ER), Small-World (SW), and Scale-Free (SF), to identify their strengths and limitations. Additionally, we discuss the implications of our findings for policy planning, highlighting how sampling methods influence the accuracy of disease spread estimates and, consequently, the effectiveness of mitigation strategies.
The extent and intensity of disease spread are driven broadly by three factors: the transmissibility of the virus, the recovery rate of individuals, and the contact network within the population (Nelson and Williams, 2014). Mechanistic models have been developed based on various combinations of these factors to study disease dynamics, predict disease trajectories, and examine the sensitivity of parameters such as transmission rate and recovery rate (Boccaletti et al., 2006, Kiss et al., 2017a, b).
Contact networks act as a means for infectious disease transmission. In a contact network, individuals are represented as nodes and an edge represents interaction between two nodes. The number of edges (connections) attached to a node is its degree (Newman, 2018). In this work, we use undirected contact networks, i.e., disease transmission can occur in both directions along an edge.
We consider three types of contact networks: 1) ER networks (Erdős and Rényi, 1959), which serve as a simple yet useful benchmark for understanding disease spread in random, unstructured networks like sexual contact networks (Holme, 2013); 2) SW networks (Watts and Strogatz, 1998), which closely mimic social mixing networks due to the high clustering and long-range links; and 3) SF networks (Li et al., 2005), characterized by the presence of hubs with high connectivity, reflecting the presence of super-spreaders in a real-world contact network (Lieberthal and Gardner, 2021). SF networks capture the extreme heterogeneity of connections found in real-world contact networks, which is less pronounced in the SW and ER networks, that have relatively uniform connectivity. Although these are theoretical networks they capture many properties of real-world networks (Aldous and Wilson, 2003). Thus, we use them here to demonstrate the diversity of behaviours that can arise.
Since our primary interest is to understand the influence of the underlying contact network on disease transmission metrics, we consider a stochastic SIR model (Brauer, 2008) to simulate disease spread on the contact network. SIR models are the most commonly used mechanistic compartmental models for infectious disease spread (Kermack and McKendrick, 1927). SIR dynamics can be simulated at either the population level or the individual level; here we do the latter.
A SIR model at the population level consists of a deterministic system of ordinary differential equations (ODEs), which describes the rate of change in the number or proportion of individuals in three compartments: susceptible, infected, and recovered. The solution predicts key features of disease spread, such as the number of infections, the rate of transmission, and the effective reproduction number (Brauer, 2008).
The SIR dynamics at the individual level, employed here, take the form of stochastic equations describing the rate of change in the probability of a population in a certain state and can account for heterogeneous mixing within the population based on the underlying contact network (Kiss et al., 2017a).
Another approach to studying disease spread is to use statistical modelling for extracting correlation patterns from data. In statistical models, usually, only a sample of the data is required. There are two key reasons for using sample data: a) the large cost and time needed in collecting data for an entire population, and b) the high computational cost of processing large datasets. The accuracy of predictions of both statistical and mechanistic models depends on the quality of the sample; if the sample is not representative of the population then it leads to biased estimates of disease metrics.
Various biases can arise from using inappropriate sampling methods, including demographic bias (Tyrer and Heyman, 2016), geographic bias (Banerjee and Chaudhury, 2010), temporal bias (Kleinbaum et al., 1981), and size bias (Arratia et al., 2019). In this study, we focus on size bias, which occurs when nodes with higher connectivity are sampled disproportionately often. For example, in survival analysis cohort studies of cancer patients, size bias can result in individuals with longer survival times being overrepresented. If uncorrected, this skews results, leading to an overestimation of survival times (Shen et al., 2009).
For diseases which are highly transmissible via close contact, such as COVID-19, measles, chickenpox etc., the number of contacts is a critical factor in the spread of the virus and it is essential to obtain an unbiased sample that accurately reflects the population’s contact distribution. Failure to do so can lead to size bias prediction of the disease spread rate and consequently ineffective mitigation.
@@
To sample data from a contact network, traversal-based sampling (TBS) methods are widely used (Gjoka et al., 2010, Craft and Caillaud, 2011, Malmros et al., 2016, Fournet and Barrat, 2017, Cui et al., 2022). These methods begin with one or more initial nodes (called seeds) and, based on certain information about their neighbours, choose the next node to sample. The primary distinction among TBS methods lies in how the next node is selected.
Two classic and frequently used probabilistic TBS methods are the RW (Wei et al., 2004) and MHRW algorithms (Hu and Lau, 2013). The RW and MHRW, as Markov chain-based algorithms share the “memoryless” property, leading to duplicates in the sample. In this work, we also compare these sampling algorithms when we remove or keep duplicate sample points.
@@
The RW algorithm is often used in epidemiological studies of large populations (Milligan et al., 2004, Qi et al., 2023) as it is relatively cheap compared to other sampling algorithms. Specifically, for infectious diseases, the RW algorithm works well for homogeneous contact networks, where individuals have a similar number of contacts. However, real-world contact networks are often heterogeneous (Nielsen et al., 2020), that is, the individuals vary in their number of contacts. This heterogeneity can result in size-biased samples when using the RW algorithm, i.e., over-representation of highly connected individuals (high-degree nodes) (Stein et al., 2014, Li et al., 2015).
@@
In (Gjoka et al., 2010), the RW and MHRW algorithms are compared for sampling online social networks, specifically Facebook. The RW algorithm produces a size-biased sample, over-representing high-degree nodes, whereas the MHRW algorithm generates a sample with a degree distribution closely matching that of the underlying Facebook network. Similar observations are reported in (Leskovec and Faloutsos, 2006b) for various network types, including citation networks from the e-print arXiv repository (Leskovec and Faloutsos, 2006b), autonomous system networks (University of Oregon, ) (Internet router networks), affiliation networks from the astrophysics category on arXiv (Leskovec and Faloutsos, 2006b), and trust networks from Epinions.com (Richardson et al., 2003). We note, however, that similar comparative analyses of RW and MHRW in the context of epidemiological research have not been undertaken, to the best of our knowledge. Here, we address this gap by exploring the RW and MHRW sampling methods for infectious disease modelling.
@@
We examine how network structures and sampling algorithms affect estimates of disease metrics. We assume a stochastic SIR model with varying transmission and recovery rates, for three network types (ER, SW and SF). Specifically, in Section 2, we quantify the size bias in the measurement of disease metrics introduced by the RW and MHRW algorithms, and in Section 3 we compare key metrics such as epidemic size, the number of secondary infections (effective reproduction number), and the time-to-infection. Finally, in Section 4, to illustrate the practical applications of our study, we analyse cattle movement network data provided by the British Cattle Movement Service (BCMS), which has hybrid properties linked to the ER, SW and SF networks. We provide a summary and directions for future research in Section 5.
@@ 2.1 Networks
We consider contact networks (Bansal et al., 2010, Craft, 2015). A contact network is defined by $G=\{V,E\}$, where $V$ is a set of node labels representing the individuals who can carry and transmit the disease, and $E$ is a set of edges that defines whether there have been interactions between the individuals. Let $N$ be the number of nodes in $G$, also known as network size. Let $k_{v}$ denote the degree of a node $v\in V$, which is defined as the number of edges (connections) of node $v$. Based on the network structure we can define and calculate (at least numerically) $p_{k}$, the probability that a node is of degree $k$, as well as the average degree of the network $\langle k\rangle$ (Newman, 2018). We will randomly generate our network structures on the basis of these degree distributions. Namely, a number of nodes will be fixed and the connections will be chosen to satisfy a given degree distribution.
@@ 2.1.1 Network Structures
Real-world networks, such as the cattle movement network from the BCMS, which we analyze in Section 4, rarely align perfectly with theoretical network structures like ER, SW, or SF networks. Instead, they often have a mixture of features from these networks, such as clustering, and heterogeneous connectivity. To better understand the implications of these structural characteristics on disease spread, we first describe these three theoretical networks individually (Newman, 2018). This approach helps us to determine how specific network features influence disease metrics, offering a baseline for understanding the more complex structure of the cattle movement network and other real networks.
@@
Figure 1: Illustration of three network structures, generated using the NetworkX (Hagberg et al., 2008) library, with $N=50$ nodes and average degree $\langle k\rangle=5$: a) ER network, b) SW network with $p=0.5$, and c) SF network with $\alpha=3$.
@@
Figure 2: Degree distribution (histogram) and fitted theoretical degree distributions (black line plots): a) the binomial distribution given by Eq. (1) with $N=50$ and $p=0.1$ for ER networks, b) a fitted Gaussian distribution with mean $4$ and standard deviation $1.2$, applied to the degree values from generated SW networks, and c) the power-law distribution given by Eq. (2) with $\alpha=3$ for SF networks. Degree distributions are estimated using $100$ simulated networks of each type with parameters $N=50$, $\langle k\rangle=5$ for the ER network, $p=0.5$ for the SW network, and $\alpha=3$ for the SF network, generated using the NetworkX (Hagberg et al., 2008) library.
@@
    Erdős-Rényi (ER) networks are generated by considering every pair of nodes and connecting them with an edge with probability $p$, independently of other edges in the network (Erdos et al., 1960). An example of an ER network is shown in Figure 1(a). ER networks follow binomial degree distribution (Newman et al., 2001), with network size $N$ and edge creation probability $p=\langle k\rangle/(N-1)$ (Figure 1(d)), that is, given by
@@ 2.2 Sampling algorithms
We compare the RW and the MHRW sampling algorithms for sampling the ER, SW, and SF networks, as described in Section 2.1.1. Below, we describe the two sampling algorithms:
@@
    The MHRW algorithm is also a traversal-based sampling method (Hu and Lau, 2013). Sampling starts by randomly selecting a node from the network. The next node is selected from the neighbours of the initial node using the proposal-acceptance mechanism (Spencer, 2021) based on the degree of the two nodes. Suppose that the last sampled node is $v$ and the proposed next node is $w$, the probability of accepting $w$ is
@@
Since the RW and MHRW are Markov chain-based algorithms and share the “memoryless” property, some nodes may be revisited during the sampling process, leading to duplicates. The duplicates introduced by revisiting nodes reveal key structural features of the network. For example, they highlight community clustering, as random walks tend to revisit nodes within tightly connected groups (Rosvall and Bergstrom, 2008). Duplicates also emphasize high-degree nodes, required for immunization and intervention strategies (Maiya and Berger-Wolf, 2011). In biological and knowledge graphs, duplicates preserve important clusters and reflect the intensity of interactions, such as gene or drug-target relationships (Dempsey et al., 2012, Zhang et al., 2021).
2.3 SIR model
In Section 2.1, we introduced the structure and properties of the networks, we now demonstrate how to apply a stochastic SIR model on three networks (ER, SW, and SF) to compare disease metric estimates derived from samples obtained using the RW and MHRW algorithms.
Consider a population of size $N$, which is categorized into three compartments according to the infection status of the individuals during an epidemic. The three compartments are Susceptible ($S$), Infected ($I$), and Recovered ($R$). Let $\bm{X}(t)\equiv(S(t),I(t),R(t))$ be the state vector where $S(t),I(t),R(t)$ are the number of individuals in state $S,I,R$ respectively at time $t$. The number of individuals in each compartment of the SIR model changes due to two events: 1) a transmission event, and 2) a recovery event, which can be written, respectively as follows,

### v3

# Reducing Size Bias in Epidemic Network Modelling

@@ Abstract
Epidemiological models help policymakers mitigate disease spread by predicting transmission metrics based on disease dynamics and contact networks. Calibrating these models requires representative network sampling. We investigate the Random Walk (RW) and Metropolis-Hastings Random Walk (MHRW) algorithms for three network types: Erdős–Rényi (ER), Small-world (SW), and Scale-free (SF). Disease transmission is simulated using a stochastic susceptible-infected-recovered (SIR) framework. For ER and SW networks, RW overestimates infected individuals and secondary infections by $25\%$ due to size bias, favouring highly connected nodes. MHRW, though more computationally intensive, reduces size bias and provides more representative samples. For time-to-infection, both algorithms provide representative estimates. However, neither algorithm samples SF networks representatively, exhibiting significant variability. Furthermore, removing duplicate sampled nodes reduces MHRW’s accuracy across three network types. We apply both algorithms to a cattle movement network of $46,512$ farms combining ER, SW, and SF features. RW overestimates infected farms by about $100\%$ and secondary infections by over $900\%$, reflecting significant size bias, while MHRW estimates align within $1\%$ of the cattle network values. RW underestimates time-to-infection by about $40\%$, while MHRW overestimates it by $10\%$. Accuracy, again, deteriorates when duplicate nodes are removed. Our findings guide algorithm selection and intervention strategies based on network structure and disease severity; RW’s conservative estimates suit high-mortality, fast-spreading epidemics, while MHRW enables more precise interventions for slower epidemics.
@@
networks , SIR model , sampling , disease modelling , size bias , cattle network , policymaking , interventions
@@
Compared Random Walk and Metropolis-Hastings Random Walk in estimating disease metrics.
Sampled Erdős-Rényi (ER), Small-world (SW), Scale-free (SF) networks.
Explored reduction of size bias, the over-representation of highly connected nodes.
MHRW outperformed RW for ER and SW networks.
RW suits fast and/or severe epidemics needing urgent interventions.
MHRW suits slower and low-severity epidemics.
@@ 1 Introduction
During an epidemic, accurately estimating metrics such as the number of infected individuals (epidemic size), effective reproduction number (secondary infections), and time-to-infection is crucial for effective disease mitigation (Nunes et al., 2024). Epidemic modelling provides a suite of tools, frameworks, and methodologies to generate disease metric estimates (Vynnycky and White, 2010). However, estimating such metrics on networks is challenging due to the complexity of contact networks (Avraam and Hadjichrysanthou, 2025), the chosen data sampling method (Danon et al., 2011), and other factors, such as data quality. In particular, inconsistency between the sampling method and the underlying contact network can lead to biased and non-representative samples, skewing disease metric estimates.
In this work, we investigate how to reduce a specific form of sampling bias known as size bias (Arratia et al., 2019), where individuals with a higher number of contacts are disproportionately overrepresented in the sample. This bias can lead to severely incorrect estimation of disease spread metrics and ineffective interventions. We aim to reduce size bias by tailoring sampling algorithms to the epidemic contact network. Our approach focuses on statistically comparing two traversal-based sampling (TBS) methods, Random Walk (RW) (Wei et al., 2004) and Metropolis-Hastings Random Walk (MHRW) (Hu and Lau, 2013), to evaluate their influence on disease metric estimation. We opt to use a stochastic susceptible-infected-recovered (SIR) model for disease spread on three network types: Erdős-Rényi (ER), Small-world (SW), and Scale-free (SF). Additionally, we demonstrate the real-world applicability of our methodology by analysing cattle movement data from the British Cattle Movement Service (BCMS) (Personal Communication, June, 14, 2024), where size bias may affect policymaking.
@@
TBS methods are widely used for sampling networks (Gjoka et al., 2010, Craft and Caillaud, 2011, Malmros et al., 2016, Fournet and Barrat, 2017, Cui et al., 2022). These methods begin with one or more initial nodes (called seeds) and, based on specific information about their neighbours, choose the next node to sample. The primary distinction among TBS methods lies in how the next node is selected. The RW and MHRW are classic and frequently used TBS algorithms.
@@
The RW algorithm is often used in epidemiological studies of large populations (Milligan et al., 2004, Qi et al., 2023) as it is relatively computationally cheap compared to other sampling algorithms. Specifically, for infectious diseases, the RW algorithm works well for homogeneous contact networks, where individuals have a similar number of contacts. However, real-world contact networks are often heterogeneous (Nielsen et al., 2020), meaning that individuals vary in their number of contacts. This heterogeneity can result in size-biased samples when using the RW algorithm (Stein et al., 2014, Li et al., 2015).
MHRW addresses the issue of size bias by adjusting the selection probability of each node based on its number of connections, thereby reducing the likelihood of sampling highly connected individuals. This correction allows MHRW to generate samples that more accurately reflect the degree distribution of the underlying network (UN). Although the MHRW algorithm is approximately $1.5-2$ times more computationally expensive compared to the RW algorithm (Bishop and Nasrabadi, 2006), it yields more representative samples, leading to more reliable statistical estimates of dynamics on heterogeneous contact networks. We note that the RW and MHRW algorithms are based on Markov chains and share the “memoryless” property, leading to duplicates in the sample. Thus, we also compare and evaluate the effect of duplicate sample removal on the accuracy of disease metric estimates for the RW and MHRW algorithms.
The extent and intensity of disease spread are driven broadly by three factors: the virus transmissibility, the recovery rate of individuals, and the population contact network (Nelson and Williams, 2014). Contact networks act as a means for infectious disease transmission. In a contact network, individuals are represented as nodes, and an edge represents the interaction between two nodes. The number of edges (connections) attached to a node is its degree (Newman, 2018). In this work, we use undirected contact networks (Newman, 2018), i.e., disease transmission can occur in both directions along an edge.
We consider three types of contact networks:1) ER networks (Erdős and Rényi, 1959), which serve as a simple yet useful benchmark for understanding disease spread in random, unstructured networks, for example, sexual contact networks (Holme, 2013); 2) SW networks (Watts and Strogatz, 1998), which closely mimic social mixing networks, with high clustering and long-range links; and 3) SF networks (Li et al., 2005), characterised by the presence of hubs with high connectivity, reflecting the presence of super-spreaders in a real-world contact network (Lieberthal and Gardner, 2021). SF networks exhibit significant heterogeneity of connections as found in real-world contact networks, whereas SW and ER networks have relatively uniform connectivity. These three theoretical networks capture many properties of real-world networks (Aldous and Wilson, 2003) and the diversity of behaviours that can arise.
We employ a mechanistic, stochastic Susceptible-Infected-Recovered (SIR) model (Brauer, 2008, Kermack and McKendrick, 1927) to simulate disease spread on networks. The individual-level SIR model is formulated using stochastic equations that describe the rate of change in the probability of individuals being in a given disease state. We chose this model because it captures heterogeneous mixing patterns within the population, as determined by the contact network structure (Kiss et al., 2017a).
In modelling disease spread, only a sample of the data is required. There are two key reasons for using sampling: a) the large cost and time needed in collecting data for an entire population and b) the high computational cost of processing large datasets. The accuracy of predictions of both statistical and mechanistic models depends on the quality of the sample; if the sample is not representative of the population contact network, then it leads to biased estimates of disease metrics (Suhail et al., 2021, Joyal-Desmarais et al., 2022)
Aside from size bias, various biases can arise from using inappropriate sampling methods, including demographic bias (Tyrer and Heyman, 2016), geographic bias (Banerjee and Chaudhury, 2010), and temporal bias (Kleinbaum et al., 1981). For diseases which are highly transmissible via close contacts, such as COVID-19, measles, chickenpox, etc., the number of contacts is a critical factor in the spread of the virus, so it is essential to obtain an unbiased sample that accurately reflects the population’s contact distribution. Failure to do so can lead to size-biased predictions of the disease spread metrics and, consequently, to ineffective mitigation strategies. Size bias is also observed in other fields, such as survival analysis of cancer patients, where size bias can result in the over-representation of individuals with longer survival times. If uncorrected, size bias skews results, leading to an overestimation of survival times (Shen et al., 2009).
@@
Previous studies have shown that the RW algorithm yields size-biased samples when applied to various heterogeneous networks. In (Gjoka et al., 2010), the RW algorithm produces a size-biased sample, over-representing high-degree nodes, whereas the MHRW algorithm generates a sample with a degree distribution closely matching that of Facebook’s network. Similar observations are reported in (Leskovec and Faloutsos, 2006b) for various network types, including citation networks from the e-print arXiv repository (Leskovec and Faloutsos, 2006b), autonomous system networks (University of Oregon, 2025) (Internet router networks), affiliation networks in the Astrophysics category on arXiv (Leskovec and Faloutsos, 2006b), and trust networks from Epinions.com (Richardson et al., 2003). However, there remains a gap in the epidemiological literature: despite the widespread use of network sampling in disease modelling, the comparative performance of RW and MHRW in estimating epidemic metrics has not been systematically assessed. To our knowledge, our study is the first to do so.
@@
This paper is structured as follows. In Section 2, we quantify size bias in the estimation of disease metrics for the RW and MHRW algorithms, and in Section 3 we compare key disease metrics such as number of infected individuals, the number of secondary infections, and the time-to-infection, for the three chosen network types (ER, SW, SF). Finally, in Section 4, to illustrate the practical applications of our study, we analyse cattle movement network data provided by the British Cattle Movement Service (BCMS) (Personal Communication, June 14, 2024), which has hybrid properties related to the ER, SW and SF networks. We provide a summary and directions for future research in Section 5.
@@ 2.1 Networks
We consider contact networks (Bansal et al., 2010, Craft, 2015). A contact network is defined by $G=\{V,E\}$, where $V$ is a set of node labels representing the individuals who can carry and transmit the disease, and $E$ is a set of edges that defines whether there have been interactions between the individuals. Let $N$ be the number of nodes in $G$, also known as network size. Let $k_{v}$ denote the degree of a node $v\in V$, which is defined as the number of edges (connections) of node $v$. Based on the network structure, we can define and calculate (at least numerically) $p_{k}$, the probability that a node is of degree $k$, as well as the average degree of the network $\langle k\rangle$ (Newman, 2018). We will randomly generate our network structures on the basis of these degree distributions. Namely, a number of nodes will be fixed, and the connections will be chosen to satisfy a given degree distribution.
@@ 2.1.1 Network structures
Real-world networks, such as the cattle movement network from the BCMS, which we analyse in Section 4, rarely align perfectly with theoretical network structures like ER, SW, or SF networks. Instead, they often have a mixture of features from these networks, such as clustering and heterogeneous connectivity. To better understand the implications of these structural characteristics on disease spread, we first describe these three theoretical networks individually (Newman, 2018). This approach helps us to determine how specific network features influence disease metrics, offering a baseline for understanding the more complex structure of the cattle movement network and other real networks.
@@
Figure 1: Illustration of three network structures, generated using the NetworkX library (Hagberg et al., 2008), with $N=50$ nodes and average degree $\langle k\rangle=5$: a) ER network, b) SW network with $p=0.5$, and c) SF network with $\alpha=3$.
@@
Figure 2: Degree distribution (histogram) and fitted theoretical degree distributions (black line plots): a) the binomial distribution given by Eq. (1) with $N=50$ and $p=0.1$ for ER networks, b) a fitted Gaussian distribution with mean $4$ and standard deviation $1.2$, applied to the degree values from generated SW networks, and c) the power-law distribution given by Eq. (2) with $\alpha=3$ for SF networks. Degree distributions are estimated using $100$ simulated networks of each type with parameters $N=50$, $\langle k\rangle=5$ for the ER network, $p=0.5$ for the SW network, and $\alpha=3$ for the SF network, generated using the NetworkX library (Hagberg et al., 2008).
@@
    Erdős-Rényi (ER) networks are generated by considering every pair of nodes and connecting them with an edge with probability $p$, independently of other edges in the network (Erdős et al., 1960). An example of an ER network is shown in Figure 1(a). ER networks follow binomial degree distribution (Newman et al., 2001), with network size $N$ and edge creation probability $p=\langle k\rangle/(N-1)$ (Figure 1(d)), that is, given by
@@ 2.2 Sampling algorithms
We compare the RW and the MHRW algorithms for sampling the ER, SW, and SF networks, as described in Section 2.1.1. Below, we first describe the algorithms:
@@
    The MHRW algorithm is also a TBS method (Hu and Lau, 2013). Sampling starts by randomly selecting a node from the network. The next node is selected from the neighbours of the initial node using the proposal-acceptance mechanism (Spencer, 2021) based on the degree of the two nodes. Suppose that the last sampled node is $v$ and the proposed next node is $w$, the probability of accepting $w$ is
@@
Since the RW and MHRW are Markov chain-based algorithms and share the “memoryless” property, some nodes may be revisited during the sampling process, leading to duplicates. The duplicates introduced by revisiting nodes reveal key structural features of the network. For example, they highlight community clustering, as random walks tend to revisit nodes within tightly connected groups (Rosvall and Bergstrom, 2008). Duplicates also emphasize high-degree nodes required for immunization and intervention strategies (Maiya and Berger-Wolf, 2011). In biological and knowledge graphs, duplicates preserve important clusters and reflect the intensity of interactions, such as gene or drug-target relationships (Dempsey et al., 2012, Zhang et al., 2021).
2.3 Stochastic SIR model
In Section 2.1, we introduced the structure and properties of the networks; we now demonstrate how to apply a stochastic SIR model on three networks (ER, SW, and SF) to compare disease metric estimates derived from samples obtained using the RW and MHRW algorithms.
Consider a population of size $N$, which is categorized into three compartments according to the infection status of the individuals during an epidemic. The three compartments are Susceptible ($S$), Infected ($I$), and Recovered ($R$). Let $\bm{X}(t)\equiv(S(t),I(t),R(t))$ be the state vector where $S(t),I(t),R(t)$ are the number of individuals in state $S,I,R$ respectively at time $t$. The number of individuals in each compartment of the SIR model changes due to two events: 1) a transmission event and 2) a recovery event, which can be written, respectively, as follows,